(4.1) Consider the point A = (−1, 0, 1), B = (0, −2, 3), and C = (−4, 4, 1) to be vertices of a

triangle ∆. Evaluate all side lengths of ∆.

(4.2) Let ∆ be the triangle with vertices the points P = (3, 1, −1), Q = (2, 0, 3) and R = (1, 1, 1).

Determine whether ∆ is a right angle triangle. If it is not, explain with reason, why?


1
Expert's answer
2021-06-23T14:28:00-0400

Solution.

"(4.1)A=(-1;0;1), B=(0;-2;3), C=(-4;4;1);"

"AB=\sqrt{(0+1)^2+(-2-0)^2+(3-1)^2}=3;"

"BC=\sqrt{(-4-0)^2+(4+2)^2+(1-3)^2}=\sqrt{56}=2\sqrt{14};"

"AC=\sqrt{(-4+1)^2+(4-0)^2+(1-1)^2}=5;"

"(4.2) P=(3;1;-1), Q=(2;0;3), R=(1;1;1)"

"PQ=\sqrt{(2-3)^2+(0-1)^2+(3+1)^2}=3\sqrt{2};"

"QR=\sqrt{(1-2)^2+(1-0)^2+(1-3)^2}=\sqrt{6};"

"PR=\sqrt{(1-3)^2+(1-1)^2+(1+1)^2}=2\sqrt{2};"

"PQ^2=QR^2+PR^2;"

"18\not =6+8; 18\not=14," therefore, the triangle is not right-angled.

Answer: "(4.1) AB=3; BC=2\sqrt{14}; AC=5;"

"(4.2)" A triangle is not right-angled because Pythagoras' theorem does not hold.